OpenStudy (anonymous):

Verify tan(x+pi/2) = -cot x

OpenStudy (anonymous):

Please help i dont understand this

OpenStudy (anonymous):

This is what i have but i dont think im doing it right tan(x+pi/2)=sin(x+pi/2)/cos(x+pi/2) sin(x)+sin(pi/2)/cos(x)+cos(pi/2) sin(pi/2)=1 cos(pi/2)=0 sin(x)+1/cos(x)

OpenStudy (anonymous):

tan(x+pi/2)=sin(x+pi/2)/cos(x+pi/2) sin(x)+sin(pi/2)/cos(x)+cos(pi/2) sin(pi/2)=1 cos(pi/2)=0 sin(x)+1/cos(x) sin(x)cos(x)+cos(x)/cos^2(x) cos^2(x)=1-sin^2(x) 1-sin^2(x)=(1+sin(x))(1-sin(x)) (cos(x))(sin(x)+1)=(1-sin(x))(1-sin(x)) cos(x)/1-sin(x) ^is this close?

OpenStudy (anonymous):

\[\tan \left( x+\frac{ \pi }{ 2 } \right)=\frac{ \tan x+\tan \frac{ \pi }{ 2 } }{ 1-\tan x~\tan \frac{ \pi }{ 2 } },\] divide the numerator and denominator by \[\tan \frac{ \pi }{ 2 }~on ~R.H.S\] \[=\frac{ \frac{ \tan x }{ \tan \frac{ \pi }{ 2 } }+\frac{ \tan \frac{ \pi }{ 2 } }{ \tan \frac{ \pi }{ 2 } } }{ \frac{ 1 }{ \tan \frac{ \pi }{ 2 } }-\frac{ \tan x \tan \frac{ \pi }{ 2 } }{ \tan \frac{ \pi }{ 2 } } }\] \[=\frac{ \tan x \cot \frac{ \pi }{ 2 } +1}{ \cot \frac{ \pi }{ 2 } -\tan x}\] \[=\frac{ \tan x *0+1 }{0-\tan x }=\frac{ 1 }{ -\tan x }=-\cot x\]

OpenStudy (anonymous):

THANK YOU SO MUCH!!!!!!!!!

OpenStudy (anonymous):

yw